OpenStudy (anonymous):

Evaluate the integral! From 0 to pi/3 of 4sec x tan x dx

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/3} 4\sec x \tan x\]

OpenStudy (anonymous):

i am helpi u right noe

OpenStudy (anonymous):

Like i get the basic jist of integrals i just get stuck on finding the antiderivative

OpenStudy (anonymous):

oka

OpenStudy (anonymous):

see if you can think of a function whose derivative is \(\sec(x)\tan(x)\)

OpenStudy (anonymous):

or look in the book for the derivatives of trig functions. you will see it

OpenStudy (anonymous):

i cant find the anser

OpenStudy (anonymous):

sec 4x and they just applied the chasin rule?

OpenStudy (anonymous):

chain*

OpenStudy (anonymous):

easier than that

OpenStudy (anonymous):

if you take the derivative of \(\sec(4x)\) you will get \(4\sec(4x)\tan(4x)\) which is not at all what you want. just leave the constant alone and take \(4\sec(x)\)

OpenStudy (anonymous):

plug in \(\frac{\pi}{3}\) plug in \(0\) and subtract actually forget 0 because \(\tan(0)=0\)

OpenStudy (anonymous):

ok perfect i can do it from here thanks :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

answers 2 right?

OpenStudy (anonymous):

yes